# Say My Name: a Model’s Bias Discovery Framework

### v1

@@ Abstract
In the last few years, due to the broad applicability of deep learning to downstream tasks and end-to-end training capabilities, increasingly more concerns about potential biases to specific, non-representative patterns have been raised. Many works focusing on unsupervised debiasing usually leverage the tendency of deep models to learn “easier” samples, for example by clustering the latent space to obtain bias pseudo-labels. However, the interpretation of such pseudo-labels is not trivial, especially for a non-expert end user, as it does not provide semantic information about the bias features. To address this issue, we introduce “Say My Name” (SaMyNa), the first tool to identify biases within deep models semantically. Unlike existing methods, our approach focuses on biases learned by the model. Our text-based pipeline enhances explainability and supports debiasing efforts: applicable during either training or post-hoc validation, our method can disentangle task-related information and proposes itself as a tool to analyze biases. Evaluation on traditional benchmarks demonstrates its effectiveness in detecting biases and even disclaiming them, showcasing its broad applicability for model diagnosis.
@@ 1 Introduction
Figure 1: Our approach, SaMyNa, searches potential spurious features learned by a model, providing a ranked list of keywords.
@@
In the past decade, advances in technology have made it possible to widely use deep learning (DL) techniques, greatly impacting the computer vision field. By allowing systems to be trained end-to-end, DL offers potentially fast model deployability to solve complex problems, leading to fast progress and changing how we perceive and analyze visual information. Today, deep learning is applied in various real-world scenarios, such as self-driving cars [1], medical imaging [2], and augmented reality [3].
@@
These huge possibilities hinder potential pitfalls. One challenge to face when deploying DL solutions lies in guaranteeing that the model does not over-rely on specific patterns that are non-representative of the real-world data distribution [4, 5]. This is key for safety, fairness, and ethics [6]. To provide a simple example, when deploying solutions in autonomous driving, simple tasks like pedestrian detection might be implicitly solved by associating specific background elements (like sidewalks or pedestrian crossings) with the presence of pedestrians. This reliance on environmental cues can act as shortcuts for the network [7], leading to poor performance when the context changes, such as when a pedestrian is crossing a road without a marked crossing lane. DL models, if not discouraged, tend to rely on spurious correlations captured at training time, especially when they are easier to learn than the actual semantic attributes [8]. We refer to these spurious correlations as biases, and we say that the model that learned such shortcuts is biased to them.
In 2021, the European Commission introduced the Artificial Intelligence Act (AI Act) to regulate AI based on the potential risks it poses [9]. Like the General Data Protection Regulation (GDPR) [10], the AI Act could set in the next few years a global standard. Ensuring that DL models avoid spurious biases that could affect their safety, trust, and accountability is not only essential for user safety but might soon also become a legal requirement.
@@
A massive recent effort has been conducted by the Computer Vision community to try to discourage the presence of biases. In a nutshell, we can roughly distinguish three main research lines: (i) supervised, where the labels of the bias are provided [11, 12]; (ii) bias-tailored, where a hint on what the potential bias might be is provided prior to training, and an ad-hoc model is deployed to capture it [13, 14]; (iii) unsupervised, where biases are guessed directed within the vanilla-trained model [8, 15, 16, 17]. When deploying a DL solution in the wild, the latter line of research appears to be the best fit, as detailed information about bias is almost surely missing. Although some solutions already exist in this context [18, 8, 15, 19], there is still a gap in the literature related to the problem of naming a specific bias affecting a DL model, providing natural language descriptors that can be directly interpretable by a human. Existing solutions either start from a predefined set of attributes [20, 21] or require computationally expensive operations, as a captioning of the entire training set [18].
Our work here moves the first steps to mine specific model’s biases. Unlike approaches that discover biases in the dataset, our focus is oriented on naming the biases captured by the model under exam. We mine the specific features in common with these samples and we associate semantic (textual) meaning to them. From this, an expert user (or prospectively a certifier software) can search and discriminate whether the learned feature is a bias or rather, a feature for the system (Fig. 1). Through “Say My Name” (SaMyNa), we aim at providing a tool that enhances explainability for the DL model’s learned features, on top of which, if necessary, any state-of-the-art debiasing approach can be used to sanitize the model.
@@
  • We propose a tool able to potentially give a name to biases in DL models. We propose a pipeline where the whole process is text-based: contrarily to end-to-end approaches, all the intermediate steps of our approach are humanly readable and interpretable (Sec. 3.2).
  • Within our approach, we can focus on potential unwanted biases by disentangling specific work domains where the model is also tested by leveraging the embedding space of a text encoder (Sec. 3.2.4).
  • Our approach is either usable when training a model, or directly at inference time. For the first, we propose a simple yet effective strategy to mine biases directly on the training set (Sec. 3.1).
  • We test our approach on well-known setups, finding the biases acknowledged by the community (Sec. 4.3.1). Besides, we also test on ImageNet-A, distinguishing cases where a bias exists and where, on the contrary, the issue is not bias-related (Sec. 4.3.2). This demonstrates the broad applicability of SaMyNa even for more general DL model diagnosis.
@@ 2 Related Works
Supervised Debiasing. Supervised methods require explicit knowledge of the bias, generally in the form of labels, indicating whether a sample presents a certain bias or not. One of the most typical approaches consists of training an explicit bias classifier, trained on the same representation space as the target classifier, in an adversarial way, forcing the encoder to extract unbiased representations [22, 23]. Alternatively, bias labels can be exploited to identify pre-defined groups, training a model to minimize the worst-case training loss, thus pushing the model towards learning biased samples  [24]. Another possibility is represented by regularization terms, which aim at achieving invariance to bias features [11, 25].
@@
Bias-Tailored Debiasing. Bias-tailored approaches usually rely on some kind of knowledge about the bias nature. For example, if the bias is textural, then custom architectures can be designed to be more sensitive to textural information. For example, [26] propose ReBias, where a custom texture bias-capturing model is designed using 1x1 convolutions. A similar approach is followed by [12], where a BagNet-18 [27] is used as a bias-capturing model.
@@
Unsupervised Debiasing. Differently from the previously described approaches, unsupervised debiasing methods do not assume any prior knowledge of the bias, facing a more realistic situation where bias is unknown. Nam et al. [8] propose LfF, where a vanilla bias-capturing model is trained with a focus on easier samples (bias-aligned), using the Generalized Cross-Entropy (GCE) loss [28], while a debiased network is trained by giving more importance to the samples that the bias-capturing model struggles to discriminate. Ji et al. [19] propose an unsupervised clustering method that learns representations invariant to some unknown or “distractor” classes in the data, by employing over-clustering. A set of unsupervised methods relies on the assumption that bias-conflicting samples are likely to be misclassified by a biased model [29, 30]. In [30] a model is trained for a few epochs and then used in inference on the training set, considering misclassified samples as bias-conflicting and vice-versa. The debiasing is then performed by up-sampling the predicted bias-conflicting samples. In [29] the training set is split into a fixed number of subsets, training a model on each of them. Then, the trained models are ensembled into a bias-commmittee and the entire training set is fed to the committee, proposing that debiasing can be performed using a weighted ERM, where the weights are proportional to the number of models in the ensemble misclassifying a certain sample. Similarly to [15], we are able to identify during the training of a vanilla model in which moment the bias is potentially best fitted by the model, with the advantage of working directly at the output of the same model instead of mining the information in its latent space. This comes both with computational advantages (given that the latent space is typically higher dimensional) and with better interpretability of the outcome, given that we work in the model’s output space.
@@
Bias Naming. Recently, methods exploiting natural language and vision-language models to identify and mitigate bias have been proposed [20, 18, 31]. Zhang et al. [31] introduce a method capable of determining subsets of images with similar attributes systematically misclassified by a model (i.e., error slices) and a rectification method based on language. However, it starts from a pre-defined set of attributes, thus hindering the possibility of discovering completely unknown and multiple biases. [20] exploits a cross-modal embedding space to identify error slices, providing natural language predictions of the identified slices. Wiles et al. [21] propose a method for automatically determining a model’s failures, exploiting large-scale vision-language models and captioners to provide interpretable descriptors of such failures in natural language. In [18], the authors propose to extract class-wise keywords representative of bias, later used for model debiasing, exploiting group-DRO [24] on the identified groups. In this work, a CLIP score is defined using the similarity between extracted keywords and correctly and misclassified samples (class-wise) and used to find the keywords associated with a bias. In [32], the authors use large-language models and text prompts for bias discovery in text-to-image generative models. Differently from the previously cited works, we introduce an unsupervised method for diagnosing a model dependency on bias for image classification tasks, that can either be performed during or after training and only exploits task-related knowledge to provide a transparent analysis on the potential hidden biases captured by the model.
@@ 3 Method
In this section, we present our proposed method SaMyNa to identify potential spurious correlations learned by the model under analysis (Sec. 3.2). Our method can be plugged to perform model diagnosis either at training time or at test: for the first case, we also present an approach to identify at which point to mine a potential bias for the target model (Sec. 3.1).
@@ 3.1 Mining Model’s Biases
Consider a supervised classification setup (having $C$ target classes), where we learn from a dataset $\mathcal{D}_{\text{train}}$ containing $N$ input samples $(\boldsymbol{x}_{1},\ldots,\boldsymbol{x}_{N})\in\mathcal{X}$, each with a corresponding ground truth label $(\hat{y}_{1},\ldots,\hat{y}_{N})\in\mathcal{Y}$. A deep neural network $\mathcal{M}$, trained for $t$ iterations, produces an output distribution $\boldsymbol{y}_{t,n}\in\mathbb{R}^{C}$ over the $C$ classes, for each input $\boldsymbol{x}_{n}$ (typically, the activation of the last layer is a softmax). The network is trained to match the ground truth label $\hat{y}_{n}$ by minimizing a loss function such as cross-entropy.
@@
If any bias is present in the training set, however, the learning process could drive the model towards the selection of spurious features [24, 8, 26], resulting in misclassification errors. Recent findings show that it is possible to identify the moment when the model best fits the bias in the training set [15]. We will formulate here the problem of identifying, at training time, when the trained model maximally fits a potential bias.
@@
We say that $\mathcal{M}$ misclassifies the $n$-th sample at the $t$-th learning iteration if $\hat{y}_{n}\neq\operatorname{argmax}(\boldsymbol{y}_{t,n})$. Focusing on this example, we know that by minimizing the loss function we aim at increasing the value of the $\hat{y}_{n}$-th component $\boldsymbol{y}_{t,n}(\hat{y}_{n})$ (while decreasing all the others). Inspired by the Hinge loss function [33], we can define a per-sample distance metric telling us how far the $n$-th sample is from being correctly classified:
$d_{t,n}=\left\{\begin{array}[]{ll}\max(\boldsymbol{y}_{t,n})-\boldsymbol{y}_{t,n}(\hat{y}_{n})&\text{if}\operatorname{argmax}_{n}(\boldsymbol{y}_{t,n})\neq\hat{y}_{n}\\ 0&\text{otherwise}.\end{array}\right.$ (1)
@@
Intuitively, the higher (1) is, the most $\mathcal{M}_{t}$ is confident in misclassifying $n$. We are interested in finding the iteration $t^{*}$ such that the model most confidently misclassifies a pool of samples:
$t^{*}=\operatorname{argmax}_{t}\frac{1}{\sum_{n}\bar{\delta}[\hat{y}_{n}-\operatorname{argmax}(\boldsymbol{y}_{t,n})]}\sum_{n}d_{t,n},$ (2)
@@
where $\delta$ is the Kronecker delta and $\bar{\delta}=1-\delta$. When reaching $t^{*}$, the most informative samples are the misclassified ones: given that the model is most confident in misclassifying them, then the model has clearly learned some spurious features. Differently from prior works [15] speculating that misclassified samples embody a bias (with the goal of applying debiasing methods), our goal is to understand why these samples are misclassified, ultimately providing an end user of the system the possibility to acknowledge the presence of a bias. For the model $\mathcal{M}_{t}$ we will split the training dataset in a pool of correctly classified samples $\mathcal{D}_{\text{train}}^{\text{correct}}$ and misclassified $\mathcal{D}_{\text{train}}^{\text{misclass}}$.

### v2

@@ Abstract
Due to the broad applicability of deep learning to downstream tasks and end-to-end training capabilities in the last few years, increasingly more concerns about potential biases to specific, non-representative patterns have been raised. Many works focusing on unsupervised debiasing leverage the tendency of deep models to learn “easier” samples, for example, by clustering the latent space to obtain bias pseudo-labels. However, their interpretation is not trivial as it does not provide semantic information about the bias features. To address this issue, we introduce “Say My Name” (SaMyNa), a tool to identify semantic biases within deep models. Unlike existing methods, our approach focuses on biases learned by the model, enhancing explainability through a text-based pipeline. Applicable during either training or post-hoc validation, our method can disentangle task-related information and propose itself as a tool to analyze biases. Evaluation on typical benchmarks demonstrates its effectiveness in detecting biases and even disclaiming them. When sided with a traditional debiasing approach for bias mitigation, it can achieve state-of-the-art performance while having the advantage of associating a semantic meaning with the discovered bias.
*Equal contribution.
@@ 1 Introduction
In the past decade, advances in technology have made it possible to widely use deep learning (DL) techniques, greatly impacting the computer vision field. By allowing systems to be trained end-to-end, DL offers potentially fast model deployability to solve complex problems, leading to fast progress and changing how we perceive and analyze visual information. Today, deep learning is applied in various real-world scenarios, such as self-driving cars (Zhang et al., 2022a), medical imaging (Yan et al., 2024), and augmented reality (Liu et al., 2020).
These huge possibilities come with potential pitfalls. One challenge to face when deploying DL solutions lies in guaranteeing that the model does not over-rely on specific patterns that are non-representative of the real-world data distribution (Ming et al., 2022; Izmailov et al., 2022).
(a) Landbird.
(b) Waterbird.
Figure 1: Top: SaMyNa searches potential spurious features learned by a model, providing a ranked list of keywords. Bottom: Example of bias heatmaps (red = correlated with bias, blue = anti-correlated).
@@
This is key for safety, fairness, and ethics (Tartaglione et al., 2023). To provide a simple example, when deploying solutions in autonomous driving, simple tasks like pedestrian detection can be implicitly solved by associating specific background elements (such as sidewalks or pedestrian crossings) with the presence of pedestrians. This reliance on environmental cues can act as shortcuts for the network (Geirhos et al., 2020), leading to poor performance when the context changes, such as when a pedestrian is crossing a road without a marked crossing lane. DL models, if not discouraged, tend to rely on spurious correlations captured during training, especially when they are easier to learn than the actual semantic attributes (Nam et al., 2020). A visual example can be found at the bottom of Fig. 1, where a model trained on waterbirds, a dataset presenting a strong spurious correlation between target labels (waterbird and landbird) and background, indeed focuses on the latter for making its predictions. A description of how these heatmaps are constructed can be found in Sec. A of the Appendix. We refer to these spurious correlations as biases, and we say that the model that learned such shortcuts is biased towards them. In 2021, the European Commission introduced the Artificial Intelligence Act (AI Act) to regulate AI based on the potential risks it poses (Madiega, 2021). Like the General Data Protection Regulation (GDPR) (GDPR, 2016), the AI Act could set in the next few years a global standard. Ensuring that DL models avoid spurious biases that could affect their safety, trust, and accountability is not only essential for user safety but might soon also become a legal requirement.
@@
A massive recent effort has been conducted by the Computer Vision community to try to discourage the presence of biases. In a nutshell, we can roughly distinguish three main research lines: (i) supervised, where the labels of the bias are provided (Barbano et al., 2023; Hong & Yang, 2021); (ii) bias-tailored, where a hint on what the potential bias might be is provided prior to training, and an ad-hoc model is deployed to capture it (Bahng et al., 2020b; Wang et al., 2019); (iii) unsupervised, where biases are guessed directly within the vanilla-trained model (Nam et al., 2020; Nahon et al., 2023; Creager et al., 2021; Li et al., 2022). When deploying a DL solution in the wild, the latter line of research appears to be the best fit, as detailed information about bias is almost surely missing. Although some solutions already exist in this context (Kim et al., 2024; Nam et al., 2020; Nahon et al., 2023; Ji et al., 2019), there is still a gap in the literature related to the problem of naming a specific bias affecting a DL model, providing natural language descriptors that can be directly interpretable by a human. Existing solutions either start from a predefined set of attributes (Eyuboglu et al., 2022; Wiles et al., 2023) or explicitly require the availability of a validation set known to contain bias-conflicting samples (Kim et al., 2024), and require to perform computationally expensive operations such as captioning the entire validation set, which may be unavailable or made up of only aligned samples. Our method is effective by using only the training set, relaxing the assumptions of existing works (Kim et al., 2024), thus setting a first step towards mining specific model biases in realistic scenarios. Unlike approaches that discover biases in the dataset, our focus is on naming the biases captured by the model under examination. We mine the specific features in common with these samples and we associate semantic (textual) meaning to them. From this, an expert user (or prospectively a certifier software) can search and discriminate whether the learned feature is a bias or rather a feature for the system (Fig. 1). Through “Say My Name” (SaMyNa), we aim at providing a tool that enhances explainability for the DL model’s learned features, on top of which, if necessary, any state-of-the-art debiasing approach can be used to sanitize the model.
@@
  • We propose a human-readable text-based pipeline for discovering and naming biases in DL models, which:
    • does not require a validation set, contrarily to existing approaches (Kim et al., 2024), and remains human-readable and interpretable at every step (Sec. 3.2);
    • leverages the embedding space of a text encoder to create learned class representations that capture potential biases of the model (Sec. 3.2.4);
    • can be applied both at training and inference time, with the former relying on a simple yet effective strategy for discovering potential biases of the model directly on the training set (Sec. 3.1).
  • We validate our approach on established benchmarks, finding biases well-known by the community (Sec. 4.2.1). Furthermore, we also test on ImageNet-A, distinguishing cases where a bias exists and where, on the contrary, generalization issues are not bias-related (Sec. 4.2.2). This demonstrates the broad applicability of SaMyNa even for more general DL model diagnosis.
  • After bias discovery (Sec. 3.1) and naming (Sec. 3.2), our method, when coupled with a standard debiasing strategy, can attain debiasing results in line with the state-of-the-art (Sec. 4.3), with the big advantage of assigning semantic meaning to the discovered bias.
@@ 2 Related Works
Supervised methods require explicit knowledge of the bias, generally in the form of labels, indicating whether a sample presents a certain bias or not. One of the most typical approaches consists of training an explicit bias classifier, trained on the same representation space as the target classifier, in an adversarial way, forcing the encoder to extract unbiased representations (Alvi et al., 2018; Xie et al., 2017). Alternatively, bias labels can be exploited to identify pre-defined groups, training a model to minimize the worst-case training loss, thus pushing the model towards learning biased samples  (Sagawa* et al., 2020). Another possibility is represented by regularization terms, which aim at achieving invariance to bias features (Barbano et al., 2023; Tartaglione et al., 2021).
Bias-Tailored Debiasing.
@@
Bias-tailored approaches usually rely on some kind of knowledge about the bias nature. For example, if the bias is textural, then custom architectures can be designed to be more sensitive to textural information. For example, (Bahng et al., 2020a) propose ReBias, where a custom texture bias-capturing model is designed using 1x1 convolutions. A similar approach is followed by (Hong & Yang, 2021), where a BagNet-18 (Brendel & Bethge, 2019) is used as a bias-capturing model.
Unsupervised Debiasing.
@@
Differently from the previously described approaches, unsupervised debiasing methods do not assume any prior knowledge of the bias, facing a more realistic situation where bias is unknown. (Nam et al., 2020) propose LfF, where a vanilla bias-capturing model is trained with a focus on easier samples (bias-aligned), using the Generalized Cross-Entropy (GCE) loss (Zhang & Sabuncu, 2018), while a debiased network is trained by giving more importance to the samples that the bias-capturing model struggles to discriminate. (Ji et al., 2019) propose an unsupervised clustering method that learns representations invariant to some unknown or “distractor” classes in the data, by employing over-clustering. A set of unsupervised methods relies on the assumption that bias-conflicting samples are likely to be misclassified by a biased model (Kim et al., 2022a). In (Liu et al., 2021), a model is trained for a few epochs and then used in inference on the training set, considering misclassified samples as bias-conflicting and vice-versa. The debiasing is then performed by up-sampling the predicted bias-conflicting samples. In (Kim et al., 2022a), the training set is split into a fixed number of subsets, training a model on each of them. Then, the trained models are ensembled into a bias-commmittee and the entire training set is fed to the committee, proposing that debiasing can be performed using a weighted ERM, where the weights are proportional to the number of models in the ensemble misclassifying a certain sample. Other lines of work, such as tackling training data containing noisy labels (Pleiss et al., 2020), introduced techniques to identify problematic training samples through margin-like measures. Similarly to (Nahon et al., 2023), we are able to identify during the training of a vanilla model in which moment the bias is potentially best fitted by the model, with the advantage of working directly at the output of the same model instead of mining the information in its latent space. This comes both with computational advantages (given that the latent space is typically higher dimensional) and with better interpretability of the outcome, given that we work in the model’s output space.
Bias Naming.
@@
Recently, methods exploiting natural language and vision-language models to identify and mitigate bias have been proposed (Eyuboglu et al., 2022; Kim et al., 2024). (Zhang et al., 2023) introduce a method capable of determining subsets of images with similar attributes systematically misclassified by a model (i.e., error slices) and a rectification method based on language. However, it starts from a pre-defined set of attributes, thus hindering the possibility of discovering completely unknown and multiple biases. (Eyuboglu et al., 2022) exploits a cross-modal embedding space to identify error slices, providing natural language predictions of the identified slices. (Wiles et al., 2023) propose a method for automatically determining a model’s failures, exploiting large-scale vision-language models and captioners to provide interpretable descriptors of such failures in natural language. In (Kim et al., 2024), the authors propose to extract class-wise keywords representative of bias, later used for model debiasing, exploiting group-DRO (Sagawa* et al., 2020) on the identified groups. In their work, a CLIP score is defined using the similarity between extracted keywords and correctly and incorrectly misclassified samples (class-wise) and used to find the keywords associated with a bias. In (D’Incà et al., 2024), the authors use large-language models and text prompts for bias discovery in text-to-image generative models. An alternative framework is proposed in (Guimard et al., 2025), where bias proposals are obtained relying on an LLM fed with task descriptions, including the textual name of the classes. A text-to-image retrieval approach, based on captions generated from the identified proposals, is then exploited to obtain biased images from an external large-scale dataset, further employed to evaluate the target model. Differently from the previously cited works, we introduce an unsupervised method for diagnosing a model dependency on bias for image classification tasks, which can either be performed during or after training and only exploits task-related knowledge to provide a transparent analysis of the potential hidden biases captured by the model.
@@ 3 Method
In this section, we present our proposed method SaMyNa to identify potential spurious correlations learned by the model under analysis (Sec. 3.2). Our method can perform model diagnosis either at training or at test time: for the first case, we also present an approach to identify at which iteration to mine a potential bias for the target model (Sec. 3.1).
@@ 3.1 Mining Model’s Biases
Consider a supervised classification setup (having $C$ target classes), where we learn from a dataset $\mathcal{D}_{\text{train}}$ containing $N$ input samples $(\bm{x}_{1},\ldots,\bm{x}_{N})\in\mathcal{X}$, each with a corresponding ground truth label $(y_{1},\ldots,y_{N})\in\mathcal{Y}$. A deep neural network $\mathcal{M}$, trained for $t$ iterations, produces an output distribution $\bm{\hat{y}}_{t,n}\in\mathbb{R}^{C}$ over the $C$ classes, for each input $\bm{x}_{n}$ (typically, the activation of the last layer is a softmax). The network is trained to match the ground truth label $y_{n}$ by minimizing a loss function such as cross-entropy.
@@
If any bias is present in the training set, however, the learning process could drive the model towards the selection of spurious features Sagawa* et al. (2020); Nam et al. (2020); Bahng et al. (2020a), resulting in misclassification errors. Recent findings show that it is possible to identify the moment when the model best fits the bias in the training set (Nahon et al., 2023). We will formulate here the problem of identifying, at training time, when the trained model maximally fits a potential bias.
Figure 2: Pipeline for SaMyNa. Given a model, we can tell on either $\mathcal{D}^{\text{train}}$or $\mathcal{D}^{\text{val}}$, which are the correctly (with green border) and the incorrectly (red border) classified samples. Amongst these, we first perform a sample subset selection looking at the latent space of the model under analysis and choosing through $k$-medoids, the most representative samples for the learned class. Then, we employ a captioner to get a textual description of these samples. From these descriptions, we extract non-rare words as keywords, and, in parallel, working in the latent space of a text encoder, we extract the mean description for the learned classes, cleansed from common features within the dataset. We finally compare this representation with the embedding of the keywords, revealing learned correlations that extend beyond the target class signal.
@@
We say that $\mathcal{M}$ misclassifies the $n$-th sample at the $t$-th learning iteration if $y_{n}\neq\operatorname*{arg\,max}(\bm{\hat{y}}_{t,n})$. Focusing on this example, we know that by minimizing the loss function we aim at increasing the value of the $y_{n}$-th component of the model’s output, denoted as $\bm{\hat{y}}_{t,n}(y_{n})$, while decreasing all the others. To measure how far the model is from predicting the desired category, we can define a per-sample distance metric telling us how far the $n$-th sample is from being correctly classified:
$d_{t,n}=\left\{\begin{array}[]{ll}\max(\bm{\hat{y}}_{t,n})-\bm{\hat{y}}_{t,n}(y_{n})&\text{if}\operatorname*{arg\,max}_{n}(\bm{\hat{y}}_{t,n})\neq y_{n}\\ 0&\text{otherwise}.\end{array}\right$. (1)
To give an intuition, the distance-metric formalized in equation 1 is inspired by the Hinge Loss (Rosasco et al., 2004), and could actually be formulated directly as $\max(\bm{\hat{y}}_{t,n})-\bm{\hat{y}}_{t,n}(\hat{y}_{n})$, as we use one-hot-encoded vectors for labels. Here, for each sample, we define the target class as the category for which the model yields the highest prediction confidence, regardless of its correctness. We then compute a per-sample metric as the margin between the confidence in this target class and the confidence in the actual ground-truth class. This value quantifies the confidence gap the model must overcome to make a correct prediction. By definition, this metric is zero when a sample is correctly classified and positive otherwise.
@@
Intuitively, the higher $d_{t,n}$ is, the most $\mathcal{M}_{t}$ is confident in misclassifying $\bm{x}_{n}$ as a specific class. We are interested in finding the iteration $t^{*}$ such that the model most confidently misclassifies a pool of samples, as in equation 2,
$t^{*}=\operatorname*{arg\,max}_{t}\frac{1}{|\mathcal{D}_{\text{train}}^{\text{misclass}}|}\sum_{n}d_{t,n},$ (2)
@@
where $\mathcal{D}_{\text{train}}^{\text{misclass}}$ is the set of misclassified training samples, and $|\mathcal{D}_{\text{train}}^{\text{misclass}}|$ is its cardinality. When reaching $t^{*}$, the most informative samples are the misclassified ones: given that the model is most confident in misclassifying them, the model may have learned some spurious features deeply affecting its final predictions. Unlike prior works (Nahon et al., 2023) that speculate that misclassified samples embody a bias (with the goal of applying debiasing methods), our goal is to understand why these samples are misclassified, ultimately providing the end user of the system the possibility to acknowledge the presence of a bias. For the model $\mathcal{M}_{t}$ we will split the training dataset in a pool of correctly classified samples $\mathcal{D}_{\text{train}}^{\text{correct}}$ and misclassified $\mathcal{D}_{\text{train}}^{\text{misclass}}$. Examples of the vanilla model’s output softmax distribution during training can be found in the Appendix (Sec. C.1).